OpenStudy (anonymous):

Find the constant of variation for the quadratic variation. 6y = 4x2 A. 3/2 B. 2/3 C. –2 D. 4/9

OpenStudy (mathstudent55):

For the quadratic variation shown below, \(y = kx^2\), the constant of variation is \(k\). Solve your equation for y.

OpenStudy (anonymous):

How can you explain?

OpenStudy (mathstudent55):

Direct variation is \(y = kx\), with constant of variation \(k\). Quadratic variation is \(y = kx^2\), with constant of variation \(k\). You have quadratic variation \(6y = 4x^2\) If you solve it for y, you will have it in \(y = kx^2\) form, and the number that multiplies x is the constant of variation.

OpenStudy (anonymous):

So what do I do with the 6

OpenStudy (mathstudent55):

Since y is being multiplied by 6 and you want y by itself, divide both sides by 6.

OpenStudy (anonymous):

ISn't the other gonna be 4x^2/6

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):